Let x be the average number of employees in a group health insurance plan, and let y be the average administrative cost as a percentage of claims.

x 3 7 15 33 74
y 40 35 30 26 16

(b) Would you say the correlation is low, moderate, or strong? positive or negative?

moderate and positive strong and negative     moderate and negative low and negative strong and positive low and positive


(c) Use a calculator to verify that Σx = 132, Σx2 = 6848, Σy = 147, Σy2 = 4657, and Σxy = 2857. Compute r. (Round your answer to three decimal places.)
r =

As x increases, does the value of r imply that y should tend to increase or decrease? Explain.

Given our value of r, we cannot draw any conclusions for the behavior of y as x increases. Given our value of r, y should tend to remain constant as x increases.     Given our value of r, y should tend to increase as x increases. Given our value of r, y should tend to decrease as x increases.
Answer #1

V Console ! > X=C(3,7,15,33,74) > y=C(40,35,30,26,16) > plot(x,y,main = scatter plot,pch=20) > correlation=cor (x,y) > corrscatter plot 40 35 30 25 20 10 20 30 40 50 60 70De Here.....= -0.9642 ...Which..means...comelatian. is....Negative.... storong........Strongly...negatiue... Cs Ex.-32.. 147.